OpenStudy (anonymous):

With one die, Andre wants to see how many times it takes to roll a number less than 5. If the die has faces with 1, 2, 3, 4, 5, and 6 dots, what is the probability in simplest form he will roll a number less than 5?

OpenStudy (hoblos):

the possible numbers less than 5 are 1,2,3,4 --> 4 possibilities the total number of possibilities is 6 so P=4/6 = 2/3

OpenStudy (anonymous):

Wait, wouldn't it be 5/6?

OpenStudy (anonymous):

"Less than 5". 5 is not less than 5 so only 4/6 combinations satisfy it.